OpenStudy (luigi0210):

Stoke's Theorem:

OpenStudy (luigi0210):

Use Stoke's theorem to solve \(\large \int_{C} F~ dr \) where F(x,y,z) =<xz, 2xy, 3xy> and C is the boundary of part of plane 3x+y+z=3 in the first octant.

OpenStudy (astrophysics):

I had trouble understanding Stoke's and I've always wanted to go over it again and learn it a better way...mhm lets see if we can figure this out \[\int\limits_{C} \vec F \cdot d \vec r = \int\limits \int\limits_S curl \vec F \cdot d \vec S\] right

OpenStudy (luigi0210):

Not sure, I'm a bit confused because of these: http://prntscr.com/7t4ipd

OpenStudy (luigi0210):

Same thing?

OpenStudy (dan815):

yes all same thing

OpenStudy (astrophysics):

Yes, you should understand the notation and what it means then I think it would be more clear hmm. I'm thinking we can pick any surface S with the boundary C.

OpenStudy (astrophysics):

This would be a nice question to observe haha, I sort of have an idea, but I think @dan815 @ganeshie8 @Empty would be better in teaching this, plus I can learn from them to xD.

OpenStudy (dan815):